OpenStudy (anonymous):

the area of a rectangle is 12x^2-21x-6

OpenStudy (kc_kennylau):

Factorize the 3 out first

OpenStudy (mathmale):

@neshia3131: I'm sure many folks here on OpenStudy would be happy to help you. We do hope, and ask, that you please include the instructions when you post an expression such as this one. Would you do that now, please?

OpenStudy (anonymous):

3(4x+1)(x-2)

OpenStudy (kc_kennylau):

@giovanni_lagos: I understand that you're trying to help the post-er, but giving him the answer directly wouldn't help him grasp the concept. Therefore, we tend to guide the post-ers to the answer.

OpenStudy (anonymous):

Oh, well this ones on me

OpenStudy (anonymous):

Factor the 3 out of 12x-21x-2 3(4x^2-7x-2) Factor the quadratic 4x^2-7x-2. The coefficient of x^2 is 4 and the constant term is -2. The product of 4 and -2 is -8 The factors of -8 which sum to -7 are 1 and -8. So, 4x^2-7x-2 = 4x^2-8x+x-2= x(4x+1)-2(4x+1) Factor 4x+1 from x(4x+1)-2(4x+1) Answer is 3(4x+1)(x-2)

OpenStudy (mathmale):

@neshia3131: I'm noticing that we haven't heard from you at all since you posted this expression. Please, would you type in the instructions for this problem and ask any questions you may have after reading the above conversation. Thank you.